# How Safe Will I Be Given What I Saw? Calibrated Prediction of Safety Chances for Image-Controlled Autonomy

### v2

@@ Abstract
Autonomous robots that rely on deep neural network controllers pose critical challenges for safety prediction, especially under partial observability and distribution shift. Traditional model-based verification techniques are limited in scalability and require access to low-dimensional state models, while model-free methods often lack reliability guarantees. This paper addresses these limitations by introducing a framework for calibrated safety prediction in end-to-end vision-controlled systems, where neither the state-transition model nor the observation model is accessible. Building on the foundation of world models, we leverage variational autoencoders and recurrent predictors to forecast future latent trajectories from raw image sequences and estimate the probability of satisfying safety properties. We distinguish between monolithic and composite prediction pipelines and introduce a calibration mechanism to quantify prediction confidence. In long-horizon predictions from high-dimensional observations, the forecasted inputs to the safety evaluator can deviate significantly from the training distribution due to compounding prediction errors and changing environmental conditions, leading to miscalibrated risk estimates. To address this, we incorporate unsupervised domain adaptation to ensure robustness of safety evaluation under distribution shift in predictions without requiring manual labels. Our formulation provides theoretical calibration guarantees and supports practical evaluation across long prediction horizons. Experimental results on three benchmarks — racing car, cart pole, and donkey car — show that our UDA-equipped evaluators maintain high accuracy and substantially lower false positive rates under distribution shift. Similarly, world model–based composite predictors outperform their monolithic counterparts on long-horizon tasks, and our conformal calibration provides reliable statistical bounds.
@@ 1 Introduction
Autonomous robots increasingly rely on high‑resolution sensors (e.g., cameras, lidars) and deep‑learning architectures Grigorescu et al. (2020); Yurtsever et al. (2020); Liu et al. (2024). End‑to‑end reinforcement and imitation learning have become widespread, bypassing traditional components such as state estimation and planning Bojarski et al. (2016); Codevilla et al. (2018); Kiran et al. (2022). Although effective, these end‑to‑end systems often depend on complex, opaque models whose internal decision‑making processes are difficult to interpret Kuznietsov et al. (2024). This lack of transparency, combined with the inherent variability and unpredictability of real‑world conditions, makes it difficult to anticipate potential failures before they manifest  Hecker et al. (2018); Filos et al. (2020). Failures in autonomous driving can lead to catastrophic accidents, which erode public trust and hinder broader adoption Penmetsa et al. (2021). Robust safety prediction is therefore essential to ensure the reliability and acceptance of these systems, making it both a technical hurdle and a societal imperative.
@@
In the autonomous driving industry, vision-based frameworks such as Tesla Vision Tesla (2025) demonstrate the feasibility of end-to-end control using only camera inputs. By learning scene geometry and motion directly from raw images, such systems remove map and sensor dependencies and adapt more readily to new roads and dynamic conditions. However, this image-only paradigm amplifies the need for robust, well-calibrated safety prediction, as perception and control are fully integrated and lack explicit, low-dimensional state representations.
@@
Moreover, in safety-critical settings, quantifying uncertainty becomes essential to operate under partial observability or environmental variability. However, modern deep models often produce overconfident or miscalibrated outputs, undermining reliability Guo et al. (2017); Minderer et al. (2021). In autonomous robots, this can result in unsafe behavior or missed interventions Lütjens et al. (2018); Feng et al. (2019). Thus, to ensure safety, predictions must be accompanied by confidence estimates that accurately reflect real-world outcomes.
@@
Many existing safety prediction methods depend on low‑dimensional, physically meaningful states (e.g., poses and velocities) and hand‑crafted indicators that require extensive domain knowledge and offline analysis Ramesh et al. (2022b); Gautam et al. (2022b); Ruchkin et al. (2022); Conlon et al. (2024), making them impractical for online application to high-dimensional, image‑based controllers. Additionally, model‑based offline verification scales poorly to vision systems, as it demands detailed modeling of the perception subsystem Everett et al. (2021); Toufighi and Girolami (2024); Mitra et al. (2025). Similarly, black‐box statistical risk assessments still tend to assume low‐dimensional state representations and rich outcome labels (e.g., signal robustness) Donze and Maler (2010); Müller et al. (2022); Lindemann et al. (2023). Recent graph‐based models HDGT Jia et al. (2022), FutureNet Wang et al. (2024), and SemanticFormer Sun et al. (2024) represent the road environment as a graph whose nodes correspond to map elements (e.g., lanes, intersections) and dynamic agents, and whose edges encode spatial or temporal relationships. These models use the graph structure to perform multi‐agent trajectory prediction conditioned on HD maps, precomputed lane graphs, and multiple sensors inputs. This dependence incurs high annotation and maintenance costs and limits application to unmapped or rapidly changing environments.
@@
However, purely data-driven, image-based predictors themselves can suffer under distribution shift, leading to miscalibrated risk estimates and degraded accuracy when encountering novel environmental conditions Filos et al. (2020); Yang et al. (2024). Under these circumstances, the predicted vision trajectory cannot be accurately classified by a visual safety evaluator, where the small errors compound over time, producing visual trajectories whose distribution deviates substantially from the training data. As a result, a static evaluator trained offline often fails to maintain calibration and exhibits inflated false-positive rates Mao et al. (2024b), which directly undermines reliability in safety-critical decisions. This motivates the need for an adaptive mechanism that can align the evaluator to the evolving input distribution without manual relabeling.
@@
This article addresses the problem of reliable online safety prediction without access to a known, low‑dimensional dynamical state. Instead, our predictor relies on a rolling window of images, for example, the RGB frames from the most recent mm time steps of a race car’s run. Given this image history, our goal is to predict the probability that the car will stay on track for the next $k$ frames. We also require that this predicted probability is calibrated well; that is, the events forecast with probability $p$ should occur approximately a $p$ fraction of the time.
@@
We introduce a flexible family of learning pipelines, shown in Figure 1, for image‑based safety prediction, featuring: (i) deep world models that learn compact latent representations from raw observations and jointly model the dynamics to predict future observations; (ii) monolithic models that capture temporal evolution of latent safety‑critical features without reconstructing full observations; (iii) an adaptive UDA module that detects distribution shifts and selectively fine‑tunes the safety evaluator online to maintain calibration on shifted inputs and (iv) post‑hoc conformal calibration method that converts raw scores into trustworthy probabilities independently of the underlying architecture.
@@
We evaluate our pipeline on three benchmarks of long‑horizon safety prediction: racing car and cart pole from simulated OpenAI Gym environments and a physical deployment of camera-driven racing cars, known as Donkey Cars Viitala et al. (2021); Viitala (2021). Our results reveal the advantages of modular pipelines with attention mechanisms and show that predicting calibrated probabilities is easier than predicting binary safety labels over extended horizons.
@@
    A family of modular pipelines for safety prediction in image‑controlled autonomous robots.
@@
    A conformal post‑hoc calibration technique to equip safety estimates with statistical guarantees.
@@
    A comprehensive, realistic experimental evaluation that highlights the impact of modularity and calibration on long‑horizon safety prediction.
An earlier iteration of this work appeared in a conference paper by Mao et al. (2024b). This journal article substantially extends the conference version in three key ways:
@@ 2.1 Trajectory Prediction
A common way of predicting the system’s performance and safety is by inferring it from predicted trajectories. Classic approaches consider model-based prediction Lefèvre et al. (2014), for instance, estimating collision risk with bicycle dynamics and Kalman filtering Ammoun and Nashashibi (2009). One common approach with safety guarantees is Hamilton-Jacobi (HJ) reachability Li et al. (2021); Nakamura and Bansal (2023), which requires precomputation based on a dynamical model. Among many deep learning-based predictors Huang et al. (2022), a popular architecture is Trajectron++ Salzmann et al. (2020), which takes in high-dimensional scene graphs and outputs future trajectories for multiple agents. A conditional VAE Sohn et al. (2015) is adopted in Trajectron++ to add constraints at the decoding stage. This can predict more than one reasonable trajectory option under the encoding condition, which helps analyze the future safety under different ego-vehicle intentions. Learning-based trajectory predictions can be augmented with conformal prediction to improve their reliability Lindemann et al. (2023); Muthali et al. (2023). These techniques have also been adapted to motion planning settings, where adaptive conformal prediction provides calibrated guarantees for multi-agent collision avoidance Dixit et al. (2023). Recent work has further emphasized generalizability and interpretability as key desiderata for trajectory predictors, advocating for Bayesian latent-variable models to provide better uncertainty modeling and extrapolation across scenarios Lu et al. (2024).
In comparison, our work eschews handcrafted scene and state representations, instead using image and latent representations that are informed by safety. This trend is echoed in efforts to design self-aware and interpretable neural networks Itkina and Kochenderfer (2023) for robust trajectory prediction under uncertainty and interpretable world models using physically meaningful representations under dynamical constraints Peper et al. (2025); Mao and Ruchkin (2024).
@@ 2.2 Performance and Safety Evaluation
A variety of recent research enables robotic systems to self-evaluate their competency Basich et al. (2022). Performance metrics vary significantly and include such examples as the time to navigate to a goal location or whether a safety constraint was violated, which is the focus of our paper. For systems that have access to their low-dimensional state, such as the velocity and the distance from the goal, hand-crafted indicators have been successful in measuring performance degradation. These indicators have been referred to robot vitals Ramesh et al. (2022a), alignment checkers Gautam et al. (2022a), assumption monitors Ruchkin et al. (2022), and operator trust Conlon et al. (2022). Typically, these indicators rely on domain knowledge and careful offline analysis of the robotic system, which are difficult to respectively obtain and perform for high-dimensional sensors.
@@
Deep neural network (DNN) controllers are particularly vulnerable to distribution shift and difficult to formally analyze  Moreno-Torres et al. (2012); Huang et al. (2020). This challenge, combined with the critical need for safety and performance guarantees, has driven the development of both model-based and model-free prediction approaches. On the model-based side, closed-loop verification approaches perform reachability analysis to evaluate the safety of a DNN-controlled system Ivanov et al. (2021); Tran et al. (2022); however, when applied to vision-based systems, these techniques require detailed modeling of the vision subsystem and have limited scalability Santa Cruz and Shoukry (2022); Hsieh et al. (2022); Mitra et al. (2025). On the other hand, model-free safety predictions rely on the correlation between performance and uncertainty measures/anomaly scores, such as autoencoder reconstruction errors and distances to representative training data Yang et al. (2022).
Striking the balance between model-based and model-free approaches are black-box statistical methods for risk assessment and safety verification  Cleaveland et al. (2022); Zarei et al. (2020); Qin et al. (2022); Michelmore et al. (2020), such as uncertainty quantification with statistical guarantees in end-to-end vision-based control Michelmore et al. (2020). These methods require low-dimensional states and rich outcome labels (such as signal robustness) — an assumption that we are relaxing in this work while addressing a similar problem.
@@
In addition, safe planning and control solve a problem complementary to ours: they find actions to control a robot safely, usually with respect to a dynamical model, such as motion planning under uncertainty for uncertain systems Knuth et al. (2021); Hibbard et al. (2022); Chou and Tedrake (2023) and a growing body of work on control barrier functions Ames et al. (2019); Xiao et al. (2023). Some approaches are robust to measurement errors induced by learning-based perception Dean et al. (2021); Yang et al. (2023), but at their core, they still hinge on the utilization of low-dimensional states, a paradigm we aim to circumvent entirely. Importantly, one of our constraints is assuming the existence of an end-to-end controller (possibly implemented with the above methods), with no room for modifications.
@@
In the context of confidence calibration, the softmax scores of classification neural networks can be interpreted as probabilities for each class; however, standard training leads to miscalibrated neural networks Guo et al. (2017); Minderer et al. (2021), as measured by Brier score and Expected Calibration Error (ECE). Calibration approaches can be categorized into extrinsic (post-hoc) calibration added on top of a trained network, such as Platt and temperature scaling Platt (1999); Guo et al. (2017), isotonic regression Zadrozny and Elkan (2002), and histogram/Bayesian binning Naeini et al. (2015) — and intrinsic calibration to modify the training, such as ensembles Zhang et al. (2020), adversarial training Lee et al. (2018), and learning from hints DeVries and Taylor (2018), error distances Xing et al. (2019), or true class probabilities Corbiere et al. (2019). Similar techniques have been developed for calibrated regression Vovk et al. (2020); Marx et al. (2022). It is feasible to obtain calibration guarantees for safety chance predictions Ruchkin et al. (2022); Cleaveland et al. (2023), but it requires a low-dimensional model-based setting. While Lindemann et al. (2023) focus on conformal calibration of temporal logic satisfaction in structured, model-based systems, our work addresses a different challenge: calibrating safety chance predictions directly from image sequences in end-to-end controlled systems without access to low-dimensional state or explicit logical specifications. To the best of the authors’ knowledge, such guarantees have not been instantiated in a model-free autonomy setting.

### v3

# How Safe Will I Be Given What I Saw? Calibrated Safety Prediction for Image-Controlled Autonomy

@@ Abstract
Autonomous robots that rely on deep neural network controllers pose critical challenges for safety prediction, especially under partial observability and distribution shift. Traditional model-based verification techniques are limited in scalability and require access to low-dimensional state models, while model-free methods often lack reliability guarantees. This paper addresses these limitations by introducing a framework for calibrated safety prediction in end-to-end vision-controlled systems, where neither the state-transition model nor the observation model is accessible. Building on the foundation of world models, we leverage variational autoencoders and recurrent predictors to forecast future latent trajectories from raw image sequences and estimate the probability of satisfying safety properties. We distinguish between monolithic and composite prediction pipelines and introduce a calibration mechanism to quantify prediction confidence. In long-horizon predictions from high-dimensional observations, the forecasted inputs to the safety evaluator can deviate significantly from the training distribution due to compounding prediction errors and changing environmental conditions, leading to miscalibrated risk estimates. To address this, we incorporate unsupervised domain adaptation to ensure robustness of safety evaluation under distribution shift in predictions without requiring manual labels. Our formulation provides theoretical calibration guarantees and supports practical evaluation across long prediction horizons. Experimental results on three benchmarks — racing car, cart pole, and donkey car — show that our UDA-equipped evaluators maintain high accuracy and substantially lower false positive rates under distribution shift. Similarly, world model-based composite predictors outperform their monolithic counterparts on long-horizon tasks, and our conformal calibration provides reliable statistical bounds.
@@ 1 Introduction
Autonomous robots increasingly rely on high-resolution sensors (e.g., cameras, lidars) and deep-learning architectures [grigorescu_survey_2020, yurtsever_autonomous_2020, liu_multimodal_2024]. End-to-end reinforcement and imitation learning have become widespread, bypassing traditional components such as state estimation and planning [bojarski_end_2016, codevilla_conditional_2018, kiran_rl_survey_2021]. Although effective, these end-to-end systems often depend on complex, opaque models whose internal decision-making processes are difficult to interpret [kuznietsov_explainable_2024]. This lack of transparency, combined with the inherent variability and unpredictability of real-world conditions, makes it difficult to anticipate potential failures before they manifest [hecker_failure_2018, filos_distribution_shift_2020]. Failures in autonomous driving can lead to catastrophic accidents, which erode public trust and hinder broader adoption [penmetsa_effects_2021]. Robust safety prediction is therefore essential to ensure the reliability and acceptance of these systems, making it both a technical hurdle and a societal imperative.
@@
In the autonomous driving industry, vision-based frameworks such as Tesla Vision demonstrate the feasibility of end-to-end control using only camera inputs  [tesla_vision_2025]. By learning scene geometry and motion directly from raw images, such systems remove map and sensor dependencies and adapt more readily to new roads and dynamic conditions. However, this image-only paradigm amplifies the need for robust, well-calibrated safety prediction, as perception and control are fully integrated and lack explicit, low-dimensional state representations.
@@
Moreover, in safety-critical settings, quantifying uncertainty becomes essential to operate under partial observability or environmental variability. However, modern deep models often produce overconfident or miscalibrated outputs, undermining reliability [guo_calibration_2017, minderer_revisiting_2021]. In autonomous robots, this can result in unsafe behavior or missed interventions [lutjens_safe_2018, feng_can_2019]. Thus, to ensure safety, predictions must be accompanied by confidence estimates that accurately reflect real-world outcomes.
@@
Many existing safety prediction methods depend on low-dimensional, physically meaningful states (e.g., poses and velocities) and hand-crafted indicators that require extensive domain knowledge and offline analysis, making them impractical for online application to high-dimensional, image-based controllers  [ramesh_robot_vitals_2022, gautam_know_limits_2022, ruchkin_confidence_2022, conlon_event_triggered_self_assessment_2024]. Additionally, model-based offline verification scales poorly to vision systems, as it demands detailed modeling of the perception subsystem [everett_efficient_2021, toufighi_decision_2024, mitra_formal_2025]. Similarly, black‐box statistical risk assessments still tend to assume low‐dimensional state representations and rich outcome labels (e.g., signal robustness) [donze_robust_2010, muller_situational_risk_2022, lindemann_conformal_2023]. Recent graph‐based models HDGT, FutureNet, and SemanticFormer represent the road environment as a graph whose nodes correspond to map elements (e.g., lanes, intersections) and dynamic agents, and whose edges encode spatial or temporal relationships [li_hdgt_2022, chen_futurenet_2023, wang_semanticformer_2024]. These models use the graph structure to perform multi‐agent trajectory prediction conditioned on HD maps, precomputed lane graphs, and multiple sensors inputs. This dependence incurs high annotation and maintenance costs and limits application to unmapped or rapidly changing environments.
@@
However, purely data-driven, image-based predictors themselves can suffer under distribution shift, leading to miscalibrated risk estimates and degraded accuracy when encountering novel environmental conditions [filos_distribution_shift_2020, yang2024generalized]. Under these circumstances, the predicted vision trajectory cannot be accurately classified by a visual safety evaluator, where the small errors compound over time, producing visual trajectories whose distribution deviates substantially from the training data. As a result, a static evaluator trained offline often fails to maintain calibration and exhibits inflated false-positive rates, which directly undermines reliability in safety-critical decisions. This motivates the need for an adaptive mechanism that can align the evaluator to the evolving input distribution without manual relabeling [mao2024safe].
@@
This article addresses the problem of reliable online safety prediction without access to a known, low-dimensional dynamical state. Instead, our predictor relies on a rolling window of images, for example, the RGB frames from the most recent mm time steps of a race car’s run. Given this image history, our goal is to predict the probability that the car will stay on track for the next $k$ frames. We also require that this predicted probability is calibrated well; that is, the events forecast with probability $p$ should occur approximately a $p$ fraction of the time.
@@
We introduce a flexible family of learning pipelines, shown in Figure 1, for image-based safety prediction, featuring: (i) deep world models that learn compact latent representations from raw observations and jointly model the dynamics to predict future observations; (ii) monolithic models that capture temporal evolution of latent safety-critical features without reconstructing full observations; (iii) an adaptive UDA module that detects distribution shifts and selectively fine-tunes the safety evaluator online to maintain calibration on shifted inputs and (iv) post-hoc conformal calibration method that converts raw scores into trustworthy probabilities independently of the underlying architecture.
@@
We evaluate our pipeline on three benchmarks of long-horizon safety prediction: racing car and cart pole from simulated OpenAI Gym environments and a physical deployment of camera-driven racing cars, known as Donkey Cars [viitala_learning_2021, viitala_scale_2021]. Our results reveal the advantages of modular pipelines with attention mechanisms and show that predicting calibrated probabilities is easier than predicting binary safety labels over extended horizons.
@@
    A family of modular pipelines for safety prediction in image-controlled autonomous robots.
@@
    A conformal post-hoc calibration technique to equip safety estimates with statistical guarantees.
@@
    A comprehensive, realistic experimental evaluation that highlights the impact of modularity and calibration on long-horizon safety prediction.
An earlier iteration of this work appeared in a conference paper [mao2024safe]. This journal article substantially extends the conference version in three key ways:
@@ 2.1 Trajectory Prediction
A common way of predicting the system’s performance and safety is by inferring it from predicted trajectories. Classic approaches consider model-based prediction, for instance, estimating collision risk with bicycle dynamics and Kalman filtering [5284727, lefevre_survey_2014]. One common approach with safety guarantees is Hamilton-Jacobi (HJ) reachability, which requires precomputation based on a dynamical model [li_prediction-based_2021, nakamura2023online]. Among many deep learning-based predictors, a popular architecture is Trajectron++, which takes in high-dimensional scene graphs and outputs future trajectories for multiple agents [huang_survey_2022, salzmann_trajectron_2020]. A conditional VAE is adopted in Trajectron++ to add constraints at the decoding stage [NIPS2015_8d55a249]. This can predict more than one reasonable trajectory option under the encoding condition, which helps analyze the future safety under different ego-vehicle intentions. Learning-based trajectory predictions can be augmented with conformal prediction to improve their reliability [lindemann_conformal_2023, muthali_multi-agent_2023]. These techniques have also been adapted to motion planning settings, where adaptive conformal prediction provides calibrated guarantees for multi-agent collision avoidance [dixit2023adaptive]. Recent work has further emphasized generalizability and interpretability as key desiderata for trajectory predictors, advocating for Bayesian latent-variable models to provide better uncertainty modeling and extrapolation across scenarios [lu2024towards].
In comparison, our work eschews handcrafted scene and state representations, instead using image and latent representations that are informed by safety. This trend is echoed in efforts to design self-aware and interpretable neural networks [itkina2023interpretable] for robust trajectory prediction under uncertainty and interpretable world models using physically meaningful representations under dynamical constraints [peper2025four, mao2024towards].
@@ 2.2 Performance and Safety Evaluation
A variety of recent research enables robotic systems to self-evaluate their competency [basich_competence-aware_2022]. Performance metrics vary significantly and include such examples as the time to navigate to a goal location or whether a safety constraint was violated, which is the focus of our paper. For systems that have access to their low-dimensional state, such as the velocity and the distance from the goal, hand-crafted indicators have been successful in measuring performance degradation. These indicators have been referred to robot vitals, alignment checkers, assumption monitors, and operator trust [9834068, 9812030, ruchkin_confidence_2022, conlon_im_2022]. Typically, these indicators rely on domain knowledge and careful offline analysis of the robotic system, which are difficult to respectively obtain and perform for high-dimensional sensors.
@@
Deep neural network (DNN) controllers are particularly vulnerable to distribution shift and difficult to formally analyze [MORENOTORRES2012521, HUANG2020100270]. This challenge, combined with the critical need for safety and performance guarantees, has driven the development of both model-based and model-free prediction approaches. On the model-based side, closed-loop verification approaches perform reachability analysis to evaluate the safety of a DNN-controlled system [ivanov_verisig_2021, tran_verification_2022]; however, when applied to vision-based systems, these techniques require detailed modeling of the vision subsystem and have limited scalability [santa_cruz_nnlander-verif_2022, hsieh_verifying_2022, mitra_formal_2025]. Furthermore, these closed-loop verification methods fundamentally require an explicit low-dimensional dynamical model to perform reachability analysis; our framework operates end-to-end from image sequences to safety predictions without access to such a model, making these methods architecturally incompatible with our setting [ivanov_verisig_2021, tran_verification_2022]. On the other hand, model-free safety predictions rely on the correlation between performance and uncertainty measures/anomaly scores, such as autoencoder reconstruction errors and distances to representative training data [yang_interpretable_2022]. Striking the balance between model-based and model-free approaches are black-box statistical methods for risk assessment and safety verification [cleaveland_risk_2022, 10.1145/3365365.3382209, 9797578, michelmore_uncertainty_2020], such as uncertainty quantification with statistical guarantees in end-to-end vision-based control [michelmore_uncertainty_2020]. These methods require low-dimensional states and rich outcome labels (such as signal robustness) — an assumption that we are relaxing in this work while addressing a similar problem.
@@
In addition, safe planning and control solve a problem complementary to ours: they find actions to control a robot safely, usually with respect to a dynamical model, such as motion planning under uncertainty for uncertain systems and a growing body of work on control barrier functions [knuth_planning_2021, hibbard_safely_2022, chou_synthesizing_2023, ames_control_2019, xiao_safe_2023]. Some approaches are robust to measurement errors induced by learning-based perception, but at their core, they still hinge on the utilization of low-dimensional states, a paradigm we aim to circumvent entirely [dean_guaranteeing_2021, yang_safe_2023]. Importantly, one of our constraints is assuming the existence of an end-to-end controller (possibly implemented with the above methods), with no room for modifications.
@@
In the context of confidence calibration, the softmax scores of classification neural networks can be interpreted as probabilities for each class; however, standard training leads to miscalibrated neural networks, as measured by Brier score and Expected Calibration Error (ECE) [guo_calibration_2017, minderer_revisiting_2021]. Calibration approaches can be categorized into extrinsic (post-hoc) calibration added on top of a trained network, such as Platt and temperature scaling, isotonic regression, and histogram/Bayesian binning—and intrinsic calibration to modify the training, such as ensembles, adversarial training, and learning from hints, error distances, or true class probabilities [platt_probabilistic_1999, guo_calibration_2017, zadrozny_transforming_2002, naeini_obtaining_2015, zhang_mix-n-match_2020, lee_training_2018, devries_learning_2018, xing_distance-based_2019, corbiere_addressing_2019]. Similar techniques have been developed for calibrated regression [vovk_conformal_2020, marx_modular_2022]. It is feasible to obtain calibration guarantees for safety chance predictions, but it requires a low-dimensional model-based setting [ruchkin_confidence_2022, cleaveland_conservative_2023]. While recent work focuses on conformal calibration of temporal logic satisfaction in structured, model-based systems, our work addresses a different challenge: calibrating safety chance predictions directly from image sequences in end-to-end controlled systems without access to low-dimensional state or explicit logical specifications [lindemann_conformal_2023]. To the best of the authors’ knowledge, such guarantees have not been instantiated in a model-free autonomy setting.